# TextSquare: Scaling up Text-Centric Visual Instruction Tuning

### v1

@@ 5.3 Qualitative Analysis
As illustrated in Figure 4, TextSquare has a formidable capability to provide plausible explanations of the answers to questions in a variety of text-centric VQA scenarios. Figure 4(a) shows that TextSquare has simple arithmetic capabilities. Figure 4(b) shows the ability to understand textual content and provide approximate location in dense text. Figure 4(c) shows the comprehension of table structure and the ability to extract contextual information relevant to the question.
@@ 5.4 Ablation Study
In order to verify the effectiveness of Square-10M, we fine-tune the baseline model InternLM-Xcomposer2 on the public text-centric VQA instruction tuning dataset (consistent with Monkey’s training data). As shown in Table, TextSquare substantially outperforms Xcomposer2 (fine-tuned) on various text-centric VQA benchmarks by 7.77.77.77.7%, which corroborates that Square-10M can fully exploit MLLM’s ability in text-centric VQA scenarios and that a large amount of high-quality instruction tuning data has a major improvement in performance.
The Effect of Evaluation Step of the Square Strategy. As shown in Table 5, there is a distinct improvement in model performance after incorporating the evaluation of the generated VQA data, which verifies that the evaluation step of the Square strategy improves the quality of VQA instruction tuning data.
@@
Figure 5: The relationship between instruction tuning dataset scale, convergence loss, and model performance in text-centric VQA scenarios. Figure (a) and Figure (b) show the relationship between data scale and convergence loss, distinguished by a scaling of the horizontal coordinate of Figure (b) with log10. Figure (c) and Figure (d) show the relationship between data scale and model performance, distinguished by a scaling of the horizontal coordinate of figure (e) with log10.
@@ 6 Limitation
Although our approach achieves remarkable results in various scenarios, there are some limitations. Firstly, large-scale data requires plenty of GPUs for long-time training, which greatly increases the training consumption. Second, while the Square strategy improves the quality of synthetic data, it still cannot reach the human level.
@@ 7 Conclusion
In this paper, we present the Square strategy for constructing a high-quality text-centric instruction tuning dataset(Square-10M). Leveraging this dataset, TextSquare significantly surpasses recent open-source models and even achieves performance comparable to GPT4V across various benchmarks. Furthermore, we derive the relationship between instruction tuning dataset scale, convergence loss, and model performance in order to pave the way for constructing even much larger datasets. Our approach provides a data-centric perspective that revisits the role of instruction-tuning data in text-centric VQA, confirming that both the quantity and quality of data are crucial to model performance. We believe that there is a promising direction on how to further improve the data quantity and quality for closing the gap between open-source models and the leading ones.

### v3

@@ 5.3 Qualitative Analysis
As illustrated in Figure 4, TextSquare has a formidable capability to provide plausible explanations of the answers to questions in a variety of text-centric VQA scenarios. Figure 4(a) shows that TextSquare has simple arithmetic capabilities. Figure 4(b) shows the ability to understand textual content and provide an approximate location in dense text. Figure 4(c) shows the comprehension of table structure and the ability to extract contextual information relevant to the question.
@@ 5.4 Ablation Study
In order to verify the effectiveness of Square-10M, we fine-tune the baseline model InternLM-Xcomposer2 on the public text-centric VQA instruction tuning dataset (consistent with Monkey’s training data). As shown in the Table, TextSquare substantially outperforms Xcomposer2 (fine-tuned) on various text-centric VQA benchmarks by 7.77.77.77.7%, which corroborates that Square-10M can fully exploit MLLM’s ability in text-centric VQA scenarios and that a large amount of high-quality instruction tuning data has a major improvement in performance.
The Effect of the Evaluation Step of the Square Strategy. As shown in Table 5, there is a distinct improvement in model performance after incorporating the evaluation of the generated VQA data, which verifies that the evaluation step of the Square strategy improves the quality of VQA instruction tuning data.
@@
Figure 5: The relationship between instruction tuning dataset scale, convergence loss, and model performance in text-centric VQA scenarios. Figure (a) and Figure (b) show the relationship between data scale and convergence loss, distinguished by a scaling of the horizontal coordinate of Figure (b) with log10. Figures (c) and (d) show the relationship between data scale and model performance, distinguished by a scaling of the horizontal coordinate of figure (e) with log10.
@@ 6 Limitation
Although our approach achieves remarkable results in various scenarios, there are some limitations. Firstly, large-scale data requires plenty of GPUs for long-term training, which greatly increases the training consumption. Second, while the Square strategy improves the quality of synthetic data, it still cannot reach the human level.
@@ 7 Conclusion
In this paper, we present the Square strategy for constructing a high-quality text-centric instruction tuning dataset(Square-10M). Leveraging this dataset, TextSquare significantly surpasses recent open-source models and even achieves performance comparable to GPT4V across various benchmarks. Furthermore, we derive the relationship between instruction tuning dataset scale, convergence loss, and model performance in order to pave the way for constructing even larger datasets. Our approach provides a data-centric perspective that revisits the role of instruction-tuning data in text-centric VQA, confirming that both the quantity and quality of data are crucial to model performance. We believe that there is a promising direction on how to further improve the data quantity and quality for closing the gap between open-source models and the leading ones.